NEAR INFRARED SPECTROSCOPY OF CARBON DIOXIDE: ^{16}O^{12}C^{16}O LINE POSITIONS

\underline{CHARLES E. MILLER}, Department of Chemistry, Haverford College, Haverford, PA 19041 USA; LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109 USA.

High-resolution near infrared 4000 - 9000 cm^{−1} spectra of carbon dioxide have been recorded using the McMath-Pierce Fourier transform spectrometer at the Kitt Peak National Solar Observatory. Spectroscopic constants have been determined for 53 different vibrational states of the ^{16}O^{12}C^{16}O isotopologue, including 8 vibrational states for which laboratory spectra have not previously been reported. Calibration using the 2-0 band of CO near 4200 cm^{−1} and the (ν_{1} + ν_{3}) combination band of C_{2}H_{2} near 6500 cm^{−1} provides absolute line position accuracies of 2 × 10^{−5} cm^{−1} (RMS) for strong, isolated transitions throughout the observed range. Fits with RMS errors less than 3.8 × 10^{−5} cm^{−1} have been obtained for the 20013 ← 00001, 20012 ← 00001, and 20011 ← 00001 bands and RMS errors less than 6 × 10^{−5} cm^{−1} have been obtained for the 30014 ← 00001, 30013 ← 00001, 30012 ← 00001, and 00031 ← 00001 bands. The new line list satisfies the line position accuracies required for the next generation of CO_{2} remote sensing instruments, improves the capability of solar-viewing spectrometers to retrieve precise column CO_{2} measurements, and provides a secondary frequency standard in the NIR.^{a}

aPart of the research described in this paper was carried out by the Jet Propulsion Laboratory,California Institute of Technology, under contract with the National Aeronautics and Space Administration.